# PnLCalib: Sports Field Registration via Points and Lines Optimization

### v4

@@ III-B1 Architecture
Following [14], the keypoint and line extremities detection utilized a modified HRNetV2-w48 [49] network as the encoder’s backbone network. HRNetv2 [49] is a new family of convolutional networks that maintains high-resolution representations through the whole process resulting in semantically richer and spatially more precise representations. To improve the spatial resolution of the predicted heatmaps, we incorporated 2×2$\times$2 × upsampling and concatenated skip-connection features from the corresponding resolution of the convolution stem to fuse the features at different scales. The resulting feature maps are produced at half the spatial resolution of the input image, striking a balance between computational efficiency and spatial detail retention. Additionally, a more lightweight version of the backbone is achieved by reducing the dense representation layers sizes in the network’s last stages. The final predictions exhibit half the resolution of the original image, with softmax and sigmoid employed as the final activation functions for keypoints and lines, respectively.
@@ III-D Point and Line Refinement Module
[…]
@@ IV-A Datasets
SN-Calib Dataset: The SoccerNetV3-Calibration (SN23) dataset [8] comprises 22,816 images extracted from SoccerNet [12] videos and encompasses a broadcast-based multi-view nature, offering a broader range of camera perspectives beyond the main broadcast camera. Cioppa et al. [8] provide annotations for all segments of the soccer field, encompassing lines, conics and goal posts. For each visible segment on the court, at least two annotated positions are provided, optimally representing the segment in a polyline format. For the conics drawn on the pitch, the annotations consist of a list of points that roughly give the circle shape when connected. Additionally, Theiner et al. [46] provided manual camera view annotations for the SoccerNetv3-Calibration 2022 (SN22) test set version, allowing the creation of data subsets taking into account the camera view distribution.
WC14 Dataset: The WorldCup 2014 dataset (WC14) [23] stands as the reference benchmark for sports field registration and consists of 209 images from ten games for training and 186 images from other ten games for testing and the corresponding manually annotated homography matrices from the FIFA WorldCup 2014. Additionally, Theiner et al. [46] provides segment annotations in SN-Calib [8] format.
@@ IV-B Evaluation Metrics
Jaccard Index ($\text{JaC}_{\gamma}$): Following Magera et al. [28] benchmarking protocol, the evaluation relies on calculating the reprojection error between each annotated point and the line to which it belongs. Adopting a binary classification approach, each pitch segment is treated as a single entity. To be considered correctly detected, all points within the segment must have a reprojection error smaller than a threshold. The projection of pitch elements from densely sampled points of the soccer field 3D model yields a polyline for each segment. Therefore, a polyline representing a soccer field segment $s$ is classified as a true positive (TP) if $\forall p\in s:\min\left(d(p,\hat{s})\right)<\gamma$, being $\hat{s}$ the corresponding annotated segment and $\gamma$ the distance threshold in pixels. Otherwise, this segment is counted as a false positive (FP). Segments only present in the annotations are counted as false negatives (FN). Hence, the Jaccard Index for camera calibration, $\text{JaC}_{\gamma}$, at a threshold of $\gamma$ pixels is defined as:
@@
where it serves as a measure of calibration accuracy. We also measure the completeness rate (CR) as the number of camera parameters provided divided by the number of images with more than four semantic line annotations in the dataset. The final score (FS) as an evaluation criterion is calculated as the product of CR and $\text{JaC}_{5}$.
@@
Projection Error: The projection error was quantified as the average distance, in meters, between the projected points using the estimated homography and the corresponding GT. To achieve that, we uniformly sampled 2,500 pixels from the visible field area of the camera image and projected them onto the field to compute the distance. The standard dimensions of a soccer field are 105$\times$68 meters.
@@ IV-D1 Camera Calibration
In team sports such as soccer, the action takes place on a nearly planar field. Consequently, most methods utilize homography estimation to map all elements positioned on this plane but cannot project non-planar points such as points belonging to goal posts or crossbars. Conversely, in [46, 32, 14] make use of a 3D model of the soccer field to extract camera pose and intrinsic parameters directly. In homography-based approaches, parameter retrieval is accomplished through homography decomposition (HDecomp). We conduct a quantitative comparison of our proposed method to state-of-the-art approaches [5, 26, 46, 32] on the SN22-test-center dataset, comprising only images where the main camera center is visible (1,454 images). Furthermore, utilizing the SoccerNet annotation format for the WC14-test dataset provided by Theiner et al. [46], we conduct a comparison of our proposed method’s performance in camera parameter estimation on the WorldCup 2014 dataset distribution.
We report the statistics from [46] for the results of state-of-the-art approaches [5, 26, 46], along with results provided by [32]. As it is shown in Tables I-II for the SN22-test and WC14 datasets, respectively, our SV method outperforms state-of-the-art approaches across several metrics in both datasets. Additionally, the inclusion of the PnL module demonstrates its effectiveness by increasing the FS metric by $4.3\%$ on the SN22-test-center dataset and a $7.6\%$ increase on the WC14-test one. Minor variations in CR, compared to previous approaches, stem from our method’s requirement for a minimum number of visible keypoints for calibration and differences in the maximum allowable reprojection error to consider the parameter estimation valid. Moreover, variations in CR between our SV models are also due to the differing maximum allowable reprojection errors. Reprojection error is computed based on keypoint errors, meaning that while line-based refinement increases keypoint reprojection error, it results in improved overall calibration. Qualitative results demonstrating the effect of the PnL refinement module are presented in Fig. 4.
@@
We also evaluate our method on the entire SN23-test dataset, reporting statistics from [14], which represents the only comparable approach capable of extending calibration assessments to full multi-view scenarios. Although the calibration approach proposed by Mavrogiannis et al. [32] extends beyond the main camera, it is limited to calibrating the main, offside, and behind-the-goalpost cameras. Moreover, their method requires training a separate model for each camera location, whereas our approach generalizes keypoint and line extremity detection across the entire SoccerNet distribution using a single model. As shown in Table I for the SN23-test datasets, our MV model outperforms the existing approaches across all the presented metrics. Additionally, our PnL refinement module showcases a $2.8\%$ increase in the FS metric. The cost function’s weighting parameter is empirically set to $\alpha=0.5$ to maximize the FS metric. Qualitative results showcasing calibration results of our base pipeline on different camera angles are presented in Fig. 5. The significant decrease in performance compared to the SN22-test-center dataset is attributed to the challenges of calibrating images with camera poses that deviate substantially from the main camera, such as close-up shots, where few or no landmarks are visible, or fisheye shots from inside the goals, where extreme lens distortion makes calibration unable. Examples of these challenging camera shots are shown in Fig. 6.
In terms of inference speed, the plain MV model achieves 7 Hz by using the proposed image size and the software and hardware configurations described above. When incorporating the PnL refinement module, the inference speed decreases to an average of 4 Hz during evaluation on the SN23-test dataset. However, the speed for specific frames varies significantly, depending on the number of detected lines and the initialization quality.
@@ IV-D2 Homography Estimation
Dataset Approach $\text{IoU}_{\text{part}}$uparrow (%) $\text{IoU}_{\text{whole}}$uparrow (%) Proj. $\downarrow$ (m) Reproj. $\downarrow$
@@ IV-D4 Ablation Study on Refinement Contribution
The contribution of each field geometrical object for initial estimate refinement—namely points, lines, and the full PnL module—is analyzed in Table V. Similar behaviors are observed across the SoccerNet distributions: after obtaining the initial calibration estimate using the proposed keypoint sets, point-based refinement does not significantly alter the results. While point refinement slightly reduces reprojection error, leading to a minor increase in the CR metric, it simultaneously lowers accuracy metrics. In contrast, line-based refinement not only reduces reprojection error but also improves accuracy metrics, as reprojection error is calculated using points, and the maximum allowable reprojection error in the method affects these outcomes. Thus, line refinement generally produces more accurate calibration results. Finally, we confirm the previously noted effectiveness of the PnL module, which balances the contributions of both geometrical objects to deliver state-of-the-art results in the SoccerNet benchmarks.
@@
SN23-test 72.2 84.9 88.4 80.4 58.1
71.6 84.2 87.7 81.3 58.2
74.3 86.4 90.8 74.9 55.7
76.7 87.2 90.1 79.5 60.9
SN22-test -center 75.8 89.7 91.9 98.1 74.4
75.8 89.8 93.7 98.1 74.4
76.3 89.4 94.3 96.5 73.7
80.6 89.9 92.4 97.7 78.7
@@
77.6 89.8 93.7 100.0 77.6
76.5 90.8 94.5 99.4 76.1
85.2 94.0 96.1 100.0 85.2
TABLE V: Ablation study of points and lines contribution in the refinement module The table presents the impact of points and lines, both individually and jointly, on the calibration estimate refinement across the SN23-test, SN22-test-center, and WC14-test datasets.
@@ V Conclusion
In this paper, we introduce a novel framework for 3D sports field registration. Our proposed pipeline adopts a minimalist approach by solely utilizing the geometric properties of the soccer field. We demonstrate superior performance in 3D camera calibration on SoccerNet and WorldCup 2014 datasets compared to state-of-the-art methods, while also achieving comparable results in homography estimation on WorldCup 2014 and TS-WorldCup datasets. Additionally, we introduce a novel calibration refinement module that leverages field landmarks, such as keypoints and lines, which, when integrated with our calibration pipeline, further enhance performance in 3D camera calibration benchmarks and achieve state-of-the-art results in homography estimation. We also extend our pipeline to multi-view camera calibration, establishing the state-of-the-art benchmark for multi-view broadcast-based camera calibration in soccer. Our method exhibits promising results, highlighting the effectiveness of utilizing a robust field template without the need for further refinements. Furthermore, the refinement module showcases its effectiveness showing superior performance and improving calibration results when visible field landmarks are scarce. As long as the video frame distortions are not too harsh (i.e., fisheye camera shots) and a minimum of four keypoints are visible, the proposed sports field registration approach is shown to be effective. In future work, we plan to enhance our approach by incorporating temporal consistency between subsequent video frames, aligning better with the nature of sports video broadcasts.

### v5

@@ III-B1 Architecture
Following [16], the keypoint and line extremities detection utilized a modified HRNetV2-w48 [63] network as the encoder’s backbone network. HRNetv2 [63] is a new family of convolutional networks that maintains high-resolution representations through the whole process resulting in semantically richer and spatially more precise representations. To improve the spatial resolution of the predicted heatmaps, we incorporated $2\times$ upsampling and concatenated skip-connection features from the corresponding resolution of the convolution stem to fuse the features at different scales. The resulting feature maps are produced at half the spatial resolution of the input image, striking a balance between computational efficiency and spatial detail retention. Additionally, a more lightweight version of the backbone is achieved by reducing the dense representation layers sizes in the network’s last stages. The final predictions exhibit half the resolution of the original image, with softmax and sigmoid employed as the final activation functions for keypoints and lines, respectively.
@@ III-D Point and Line Refinement Module
$\begin{aligned} d(\mathbf{l},\mathbf{l}_{d})&=d(\mathbf{l},\bar{\mathbf{p}}_{d})+d(\mathbf{l},\bar{\mathbf{q}}_{d})\\ &=\frac{|\Delta_{y}\bar{\mathbf{p}}_{d,x}-\Delta_{x}\bar{\mathbf{p}}_{d,y}+\bar{\mathbf{q}}_{x}\bar{\mathbf{p}}_{y}-\bar{\mathbf{q}}_{y}\bar{\mathbf{p}}_{x}|}{\sqrt{\Delta_{y}^{2}+\Delta_{x}^{2}}}\\ &+\frac{|\Delta_{y}\bar{\mathbf{q}}_{d,x}-\Delta_{x}\bar{\mathbf{q}}_{d,y}+\bar{\mathbf{q}}_{x}\bar{\mathbf{p}}_{y}-\bar{\mathbf{q}}_{y}\bar{\mathbf{p}}_{x}|}{\sqrt{\Delta_{y}^{2}+\Delta_{x}^{2}}}\end{aligned},$ (5)
@@ IV-A Datasets
SN-Calib Dataset: The SoccerNetV3-Calibration (SN23) dataset [10] comprises 22,816 images extracted from SoccerNet [14] videos and encompasses a broadcast-based multi-view nature, offering a broader range of camera perspectives beyond the main broadcast camera. [10] provide annotations for all segments of the soccer field, encompassing lines, conics and goal posts. For each visible segment on the court, at least two annotated positions are provided, optimally representing the segment in a polyline format. For the conics drawn on the pitch, the annotations consist of a list of points that roughly give the circle shape when connected. Additionally, [59] provided manual camera view annotations for the SoccerNetv3-Calibration 2022 (SN22) test set version, allowing the creation of data subsets taking into account the camera view distribution.
WC14 Dataset: The WorldCup 2014 dataset (WC14) [26] stands as the reference benchmark for sports field registration and consists of 209 images from ten games for training and 186 images from other ten games for testing and the corresponding manually annotated homography matrices from the FIFA WorldCup 2014. Additionally, [59] provides segment annotations in SN-Calib [10] format.
@@ IV-B Evaluation Metrics
Jaccard Index ($\text{JaC}_{\gamma}$): Following [38] benchmarking protocol, the evaluation relies on calculating the reprojection error between each annotated point and the line to which it belongs. Adopting a binary classification approach, each pitch segment is treated as a single entity. To be considered correctly detected, all points within the segment must have a reprojection error smaller than a threshold. The projection of pitch elements from densely sampled points of the soccer field 3D model yields a polyline for each segment. Therefore, a polyline representing a soccer field segment $s$ is classified as a true positive (TP) if $\forall p\in s:\min\left(d(p,\hat{s})\right)<\gamma$, being $\hat{s}$ the corresponding annotated segment and $\gamma$ the distance threshold in pixels. Otherwise, this segment is counted as a false positive (FP). Segments only present in the annotations are counted as false negatives (FN). Hence, the Jaccard Index for camera calibration, $\text{JaC}_{\gamma}$, at a threshold of $\gamma$ pixels is defined as:
@@
where it serves as a measure of calibration accuracy. We also measure the completeness rate (CR) as the number of camera parameters provided divided by the number of images with more than four semantically labeled lines in the dataset. The final score (FS) as an evaluation criterion is calculated as the product of CR and $\text{JaC}_{5}$.
@@
[…]
@@ IV-D1 Camera Calibration
In team sports such as soccer, the action takes place on a nearly planar field. Consequently, most methods utilize homography estimation to map all elements positioned on this plane but cannot project non-planar points such as points belonging to goal posts or crossbars. Conversely, in [59, 42, 16] make use of a 3D model of the soccer field to extract camera pose and intrinsic parameters directly. In homography-based approaches, parameter retrieval is accomplished through homography decomposition (HDecomp). We conduct a quantitative comparison of our proposed method to state-of-the-art approaches [6, 31, 59, 42] on the SN22-test-center dataset, comprising only images where the main camera center is visible (1,454 images). Furthermore, utilizing the SoccerNet annotation format for the WC14-test dataset provided by [59], we conduct a comparison of our proposed method’s performance in camera parameter estimation on the WorldCup 2014 dataset distribution.
We report the statistics from [59] for the results of state-of-the-art approaches [6, 31, 59], along with results provided by [42]. As it is shown in Tables I-II for the SN22-test and WC14 datasets, respectively, our SV method outperforms state-of-the-art approaches across several metrics in both datasets. Additionally, the inclusion of the PnL module demonstrates its effectiveness by increasing the FS metric by $5.7\%$ on the SN22-test-center dataset and a $8.3\%$ increase on the WC14-test one. Minor variations in CR, compared to previous approaches, stem from our method’s requirement for a minimum number of visible keypoints for calibration and differences in the maximum allowable reprojection error to consider the parameter estimation valid. Moreover, variations in CR between our SV models are also due to the differing maximum allowable reprojection errors. Reprojection error is computed based on keypoint errors, meaning that while line-based refinement increases keypoint reprojection error, it results in improved overall calibration. Qualitative results demonstrating the effect of the PnL refinement module are presented in Fig. 4.
@@
We also evaluate our method on the entire SN23-test dataset, reporting statistics from [16], which represents the only comparable approach capable of extending calibration assessments to full multi-view scenarios. Although the calibration approach proposed by [42] extends beyond the main camera, it is limited to calibrating the main, offside, and behind-the-goalpost cameras. Moreover, their method requires training a separate model for each camera location, whereas our approach generalizes keypoint and line extremity detection across the entire SoccerNet distribution using a single model. As shown in Table I for the SN23-test datasets, our MV model outperforms the existing approaches across all the presented metrics. Additionally, our PnL refinement module showcases a $3.2\%$ increase in the FS metric. The weighting parameter $\alpha$ in Eq. (7) is empirically set to $\alpha=0.6$ to maximize the FS metric across camera calibration experiments. An ablation study of the calibration results with respect to different values of $\alpha$ is presented in Table V, evaluated on the SN23-test dataset. Qualitative results showcasing calibration results of our base pipeline on different camera angles are presented in Fig. 5. The significant decrease in performance compared to the SN22-test-center dataset is attributed to the challenges of calibrating images with camera poses that deviate substantially from the main camera, such as close-up shots, where few or no landmarks are visible, or fisheye shots from inside the goals, where extreme lens distortion makes calibration unable. Examples of these challenging camera shots are shown in Fig. 6.
In terms of inference speed, an ablation study on the frame rate is presented in Table VI under various configurations of heuristic voting, evaluated on the SN23-test set. Specifically, we assess different RANSAC configurations (ranging from no filtering to a $50$ px threshold) and different keypoint subsets (i.e., full-keypoints, main-keypoints, and ground-plane-keypoints). Using a fixed RANSAC threshold of $5$ px and the full-keypoints subset, our method initially achieves exhibits near-competitive performance at $164$ ms, with the PnL optimization module accounting for approximately an $8\%$ increase in inference time. By evaluating six different RANSAC configurations with a single keypoint subset, the method achieves competitive calibration performance at $273$ ms. State-of-the-art calibration accuracy is obtained when employing the full set of heuristic voting configurations, resulting in an inference time of 439 ms. All reported timings are averaged over 10 independent runs, using the proposed image resolution and the software and hardware settings described above. It is also important to note that inference time can vary significantly across frames, primarily due to differences in the number of detected lines and the quality of initialization.
@@ IV-D2 Homography Estimation
Dataset Approach $\text{IoU}_{\text{part}}$ $\uparrow$ (%) $\text{IoU}_{\text{whole}}$ $\uparrow$ (%) Proj. $\downarrow$ (m) Reproj. $\downarrow$
@@ IV-D3 Ablation Study on Keypoint Sets Contribution
$\text{JaC}_{\gamma}$ [%]
$\alpha$ 5 10 20 CR FS
0.3 73.8 86.0 89.3 80.4 59.3
0.4 75.5 86.9 89.9 79.7 60.2
0.5 76.6 87.6 90.3 79.3 60.7
0.6 78.8 89.6 91.9 78.4 61.8
0.7 78.0 88.9 91.4 78.2 61.0
TABLE V: Ablation study of the $\alpha$ parameter in the PnL optimization module, evaluated on the SN23-test dataset.
$\text{RANSAC configs}$ $\text{Keypoint subsets}$ PnL module $\text{Frame rate (ms)}\downarrow$ CR $\text{FS}\uparrow$
1 1 164 74.6 52.7
1 1 178 72.3 54.1
6 1 273 74.5 57.7
6 3 439 78.4 61.8
TABLE VI: Ablation study on the inference time of our method. The table presents the method’s frame rate under various heuristic voting configurations and with or without the use of the PnL module, evaluated on the SN23-test dataset.
@@ IV-D4 Ablation Study on Refinement Contribution
The contribution of each field geometrical object for initial estimate refinement—namely points, lines, and the full PnL module—is analyzed in Table VII. Similar behaviors are observed across the SoccerNet distributions: after obtaining the initial calibration estimate using the proposed keypoint sets, point-based refinement does not significantly alter the results. While point refinement slightly reduces reprojection error, leading to a minor increase in the CR metric, it simultaneously lowers accuracy metrics. In contrast, line-based refinement not only reduces reprojection error but also improves accuracy metrics, as reprojection error is calculated using points, and the maximum allowable reprojection error in the method affects these outcomes. Thus, line refinement generally produces more accurate calibration results. Finally, we confirm the previously noted effectiveness of the PnL module, which balances the contributions of both geometrical objects to deliver state-of-the-art results in the SoccerNet benchmarks.
@@
SN23-test 73.1 86.1 89.6 80.1 58.6
71.6 84.4 88.1 81.5 58.4
75.8 88.0 92.3 72.4 54.9
78.7 89.6 91.9 78.4 61.8
SN22-test -center 74.4 88.8 92.6 99.1 73.8
74.4 88.9 92.7 99.1 73.8
75.8 88.8 93.5 96.8 73.3
80.6 91.6 93.7 98.7 79.5
@@
77.8 89.6 93.3 100.0 77.8
77.3 91.3 95.1 96.7 74.8
85.9 93.7 95.8 100.0 85.9
TABLE VII: Ablation study of points and lines contribution in the refinement module. The table presents the impact of points and lines, both individually and jointly, on the calibration estimate refinement across the SN23-test, SN22-test-center and WC14-test datasets.
@@ V Conclusion
In this paper, we introduce a novel open-source framework for 3D sports field registration. Our proposed pipeline adopts a minimalist approach by solely utilizing the geometric properties of the soccer field. We demonstrate superior performance in 3D camera calibration on SoccerNet and WorldCup 2014 datasets compared to state-of-the-art methods, while also achieving comparable results in homography estimation on WorldCup 2014 and TS-WorldCup datasets. Additionally, we introduce a novel calibration refinement module that leverages field landmarks, such as keypoints and lines, which, when integrated with our calibration pipeline, further enhance performance in 3D camera calibration benchmarks and achieve state-of-the-art results in homography estimation. We also extend our pipeline to multi-view camera calibration, establishing the state-of-the-art benchmark for multi-view broadcast-based camera calibration in soccer. Our method exhibits promising results, highlighting the effectiveness of utilizing a robust field template without the need for further refinements. Furthermore, the refinement module showcases its effectiveness showing superior performance and improving calibration results when visible field landmarks are scarce. As long as the video frame distortions are not too harsh (i.e., fisheye camera shots) and a minimum of four keypoints are visible, the proposed sports field registration approach is shown to be effective. In future work, we plan to enhance our approach by incorporating temporal consistency across consecutive video frames, better aligning with the nature of sports video broadcasts. Additionally, a potential extension of this work includes integrating distortion modeling into the camera calibration pipeline, enabling support for a broader range of lenses—from minor distortions to fisheye effects.
Appendix A Analytical Validation of Tangent Points
Consider an estimated homography matrix $\mathbf{H}\in\mathbb{R}^{3\times 3}$, a conic represented in image coordinates by the symmetric matrix $\mathbf{\bar{c}}$, and an external point $\mathbf{\bar{p}}$. We compute the tangent lines from the external point $\mathbf{\bar{p}}$ to the conic $\mathbf{\bar{c}}$ in image coordinates. The corresponding contact points with the conic are denoted by $\mathbf{\bar{x}}$. Then, the point $\mathbf{\bar{x}}$ satisfies the following conditions:
  • $\mathbf{\bar{x}}^{\top}\mathbf{\bar{c}}\mathbf{\bar{x}}=0$, i.e., the point $\mathbf{\bar{x}}$ lies on the conic $\mathbf{\bar{c}}$ in image coordinates.
  • $\mathbf{\bar{p}}^{\top}\mathbf{\bar{c}}\mathbf{\bar{x}}=0$, i.e., the tangent line to $\mathbf{\bar{c}}$ at $\mathbf{\bar{x}}$ passes through $\mathbf{\bar{p}}$.
The conic can be reprojected into world coordinates using the inverse homography as:
$\mathbf{c}=\mathbf{H}^{-\top}\mathbf{\bar{c}}\mathbf{H}^{-1}.$
We also define the world-coordinate representations of the point $\mathbf{\bar{p}}$ and the tangent contact point $\mathbf{\bar{x}}$ as:
$\mathbf{p}=\mathbf{H}^{-1}\mathbf{\bar{p}},\quad\mathbf{x}=\mathbf{H}^{-1}\mathbf{\bar{x}}.$
Proof that $\mathbf{x}$ lies on the world conic $\mathbf{c}$
We can show that $\mathbf{x}$ lies on $\mathbf{c}$ as follows:
$\mathbf{x}^{\top}\mathbf{c}\mathbf{x}$ $=(\mathbf{H}^{-1}\mathbf{\bar{x}})^{\top}\mathbf{c}(\mathbf{H}^{-1}\mathbf{\bar{x}})$
$=\mathbf{\bar{x}}^{\top}\mathbf{H}^{-\top}\mathbf{c}\mathbf{H}^{-1}\mathbf{\bar{x}}$
$=\mathbf{\bar{x}}^{\top}\mathbf{\bar{c}}\mathbf{\bar{x}}=0.$
Hence, if $\mathbf{x}^{\top}\mathbf{c}\mathbf{x}=0$, the point $\mathbf{x}$ lies on the conic $\mathbf{c}$.
Proof that line $\mathbf{p}\mathbf{x}$ is tangent to $\mathbf{c}$ at $\mathbf{x}$
We can also show that the line joining $\mathbf{p}$ and $\mathbf{x}$ is tangent to the conic $\mathbf{c}$ at $\mathbf{x}$:
$\mathbf{p}^{\top}\mathbf{c}\mathbf{x}$ $=(\mathbf{H}^{-1}\mathbf{\bar{p}})^{\top}\mathbf{c}(\mathbf{H}^{-1}\mathbf{\bar{x}})$
$=\mathbf{\bar{p}}^{\top}\mathbf{H}^{-\top}\mathbf{c}\mathbf{H}^{-1}\mathbf{\bar{x}}$
$=\mathbf{\bar{p}}^{\top}\mathbf{\bar{c}}\mathbf{\bar{x}}=0.$
Therefore, if $\mathbf{p}^{\top}\mathbf{c}\mathbf{x}=0$, the line $\mathbf{p}\mathbf{x}$ is tangent to the conic $\mathbf{c}$ at the point $\mathbf{x}$.